# From Single to Multi: How LLMs Hallucinate in Multi-Document Summarization

### v1

@@ Abstract
Although many studies have investigated and reduced hallucinations in large language models (LLMs) for single-document tasks, research on hallucination in multi-document summarization (MDS) tasks remains largely unexplored. Specifically, it is unclear how the challenges arising from handling multiple documents (e.g., repetition and diversity of information) affect models outputs. In this work, we investigate how hallucinations manifest in LLMs when summarizing topic-specific information from multiple documents. Since no benchmarks exist for investigating hallucinations in MDS, we use existing news and conversation datasets, annotated with topic-specific insights, to create two novel multi-document benchmarks. When evaluating 5 LLMs on our benchmarks, we observe that on average, up to 75% of the content in LLM-generated summary is hallucinated, with hallucinations more likely to occur towards the end of the summaries. Moreover, when summarizing non-existent topic-related information, gpt-3.5-turbo and GPT-4o still generate summaries about 79.35% and 44% of the time, raising concerns about their tendency to fabricate content. To understand the characteristics of these hallucinations, we manually evaluate 700+ insights and find that most errors stem from either failing to follow instructions or producing overly generic insights. Motivated by these observations, we investigate the efficacy of simple post-hoc baselines in mitigating hallucinations but find them only moderately effective. Our results underscore the need for more effective approaches to systematically mitigate hallucinations in MDS. We release our dataset and code111github.com/megagonlabs/Hallucination_MDS.
@@
Catarina G. Belemthanks:  This work was conducted while Catarina G. Belem was a research intern at Megagon Labs. University of California Irvine cbelem@uci.edu                        Pouya Pezeskhpour Megagon Labs pouya@megagon.ai                        Hayate Iso Megagon Labs hayate@megagon.ai
@@ 1 Introduction
Multi-document summarization (MDS) has numerous real-world applications, including planning treatments and diagnosing patients based on their medical history (i.e., doctor notes, lab reports) (Tang et al., 2023b), forming legal arguments by linking precedents (Rodgers et al., 2023; Wu et al., 2023), or screening resumes to match candidates to job descriptions (Wang et al., 2024; Du et al., 2024). These tasks often involve linking information across lengthy documents, making the summarization process time- and labor-intensive (Van Veen et al., 2024). Recently, large language models (LLMs) have been proposed as more efficient approaches to reduce the human effort and increase scalability (Katz et al., 2023; Van Veen et al., 2024; Liu et al., 2024a).
@@
However, despite the remarkable advances of LLMs in various tasks (Bubeck et al., 2023; Zhao et al., 2024), the frequent generation of ungrounded yet plausible-sounding text, referred to as “hallucinations”, undermines trust in their output (Maynez et al., 2020; Uluoglakci and Temizel, 2024; Kalai and Vempala, 2024). While LLM hallucinations have been extensively studied in single-document tasks (Ji et al., 2023; Huang et al., 2023), resulting in the development of various evaluation benchmarks (Lin et al., 2022; Wang et al., 2022; Yang et al., 2023) and taxonomies (Rawte et al., 2023; Zhang et al., 2023; Mishra et al., 2024; Dahl et al., 2024), little is known about how processing multiple documents affects the hallucinatory behavior of LLMs in MDS. Specifically, by focusing on the aggregated quality metrics of LLM outputs, including coverage (Fabbri et al., 2019; Lu et al., 2020; Laban et al., 2024) and faithfulness (Pu et al., 2023; Huang et al., 2024), existing work offers limited insight into how the inherent challenges of MDS (e.g., repeated and diversity of information) correlates with LLM hallucinations. Consider the motivating example in Figure 1: when summarizing the repeated information from multiple documents, the LLM fails to link overlapping information across documents and satisfy the subtopic condition (highlighted text in the figure).
In this work, we comprehensively investigate the hallucinatory behavior of 5 popular LLMs in MDS tasks. Focusing on news articles and dialogues summarization, we examine the properties of hallucinations produced by LLMs when summarizing topic-specific information from multiple documents. Specifically, we aim to understand the patterns of hallucinations, their frequency, and their relationship with the number of input documents, as well as how these patterns change with the number of input documents and task focus (e.g., summarizing undiscussed topics, or focusing the summary on repeated information). To this end, we design an evaluation protocol based on fine-grained annotations concerning relevant insights—units of information—within each document. We use the insight-level annotations from the SummHay datasets (Laban et al., 2024), designed for evaluating LLMs in long-context summarization, and create benchmarks with combinations of up to 10 documents in both the conversational and news domains. Our goal here is to leverage the insight-level annotations to automatically assess the correctness of the LLM-generated summaries.
@@ 2.1 Problem Formulation
We frame the MDS task as follows: given $N$ documents ($d_{1},\dots,d_{N}$), and $K$ conditions ($c_{1},...,c_{K}$), an LLM must generate a summary $\hat{\mathbf{y}}$ such that it satisfies all $K$ conditions and is grounded in the documents. Examples of conditions include matching specific topics, adhering to length constraints, or following particular writing styles. For simplicity, we assume that the $N$ documents in our formulation are of the same type (e.g., all documents are news articles), rather than from diverse sources (e.g., job descriptions and resumes, doctor notes and prescriptions). When summarizing multiple documents, the focus can be on capturing diverse information (Huang et al., 2024) or highlighting shared commonalities. The latter is especially useful for combating misinformation, as it emphasizes “trustworthy” information corroborated by multiple sources. To analyze model behavior in both contexts, we require fine-grained annotations of the information units contained in each document.
@@ 2.2 Dataset Creation
Recently, Laban et al. (2024) introduced two MDS datasets with insight-level annotations. [^2] The authors define insights as “units of information”. The datasets cover two distinct domains—news and conversation—each containing 500 documents, evenly distributed across 5 topics, with documents averaging around 750 words. In particular, the news dataset (SummHay-News) is entity-centric and quantitative, featuring brands (e.g., Tesla), banks (e.g., JP Morgan), celebrities (e.g., Elon Musk), along with numbers and dates. In contrast, the conversation dataset (SummHay-Conv) involves everyday scenarios with 2 to 4 participants, such as medical appointments and debates. Originally developed to evaluate LLMs in long-context tasks (e.g., retriever-augmented generation, multi-document summarization), the documents were designed to ensure insights are repeated across at least 6 per topic and categorized into various subtopics (See Appendix A for more details on the SummHay).
By selecting SummHay-News and SummHay-Conv as testbeds, we can exploit their insight-level annotations to systematically assess model behavior across different settings, such as number of input documents and task focus. To control for input length (Li et al., 2024) while evaluating model behavior, we organize documents in the SummHay dataset into sets of $N$ documents, referred to as combinations of $N$ or $N$-documents (see Figure 6 in Appendix for an illustration of our benchmark creation process). Specifically, given a corpus with insight-level annotations and a subtopic $q$, we create our benchmarks by grouping documents into combinations of $N$ where multiple subtopic-related insights co-occur in two or more documents. Because we have the insight-level annotations, we can automatically identify the ground truth insights related to subtopic $q$ (called reference insights). To fully benefit from these fine-grained annotations, we instruct LLMs to succintly333 Unlike Laban et al., we do not instruct LLMs to limit their generation to a fixed number of bullet-points. Instead, we study their behavior “in the wild” when the exact number of insights is unknown. For further details related to prompts, refer to Appendix B. summarize the information as individual insights, presented in the form of bullet-point lists. We coin this setting “subtopic”. Additionally, we introduce a “subtopic+shared” setting by refining the prompt and limiting the reference insights to only the shared insights.
@@ 2.3 Automatic Evaluation
With reference insights for each combination, we can automatically evaluate the correctness of LLM outputs, distinguishing between hallucinatory and non-hallucinatory content. The correctness metric should assess whether predicted insights capture all essential information from the reference while avoiding the inclusion of irrelevant details (Huang et al., 2024). This ensures that predicted insights do not contain unsupported or potentially fabricated information.
Previous work uses an LLM to assess whether reference insights are fully, partially, or not covered by the predicted insights, showing strong correlation with human evaluations (Laban et al., 2024). However, this metric only measures reference coverage and ignores the validity of additional details in predictions. To enhance this, we assess how well predicted insights align with reference insights by applying the original metric twice—once for reference coverage and once for predicted insight—–yielding two labels per pair. We then combine these into a single correctness label using a conservative approach based on the order: NO $\prec$ PARTIAL $\prec$ FULL. By selecting the worst label, we ensure the final assessment captures discrepancies related to missing relevant information or any potentially unfaithful content that is generated by the LLM.
@@ 2.4 Metrics
A key question in MDS tasks is whether LLM-generated summaries include all correct information. To estimate average LLM correctness, we calculate macro-recall by determining the fraction of reference insights covered in each generated summary and averaging these scores across all to obtain a single score. We also report the average fraction of hallucinated content in generated summaries by computing the average proportion of predicted insights that do not correspond to any reference insight across all summaries. This metric is referred to as the false discovery rate (macro-FDR). Both metrics range from 0 to 1, with a perfect LLM scoring 1 for macro-recall and 0 for macro-FDR. For simplicity, we refer to macro-recall as recall and macro-FDR as hallucination rate.
@@ 3 Experimental Settings
Automatic Evaluation. As previously mentioned, this paper uses an LLM-as-a-judge approach (Zheng et al., 2023) similar to that of Laban et al. (2024). However, instead of using gpt-4o, we resort to a competitive yet more budget-friendly option—gpt-4o-mini-2024-07-18. The authors manually annotated 100 insight pairs and found strong alignment between both models, confirming the suitability of gpt-4o-mini-2024-07-18 as an evaluator. Refer to Appendix D for more details.
@@ 4 Experimental Results
In this section, we first investigate the extent of LLMs hallucinations in MDS under the proposed “subtopic” and “subtopic+shared” settings. Then, we analyze the models’ hallucinatory behavior when synthesizing information on non-existent topics and investigate the link between hallucinated content and characteristics of both the input and output (we use the prompts listed in Appendix B).
@@
Figure 3: Performance metrics as a function of the number of input documents in the “subtopic+shared” setting. Each line represents the mean value, with shaded areas indicating the 95% confidence intervals.
@@ 4.2 Shared Insights, Greater Errors
Compared to the previous “subtopic” setting, analyzing model behavior in “subtopic+shared” requires adjusting the evaluation process and summarization prompt to focus only on shared insights. As before, we prompt models to summarize all 500 examples per benchmark and measure recall and hallucination rates. Our findings are as follows:
LLMs generate shorter summaries but hallucinate more in general. To understand how responsive models are to the shared instruction, we compare summary lengths between “subtopic” and “subtopic+shared” settings (see Appendix E.2). Summaries in “subtopic+shared” are shorter, suggesting models respond to the shared instruction. However, shorter summaries do not guarantee quality: hallucination rates are, on average, higher in “subtopic+shared” than “subtopic” (news: +10.47%, conv: +4.20%), indicating models may struggle to identify shared insights. We also observe that, LLMs struggle to identify subtopic-related shared insights. Compared to the “subtopic” scenario, we observe an average increase in summarization performance in news domain (+6.93%) but a slight decrease in the conversation (-2.91%). Overall recall trends remain the same: larger combination sizes lead to up to a 33% drop, with the sharpest declines in GPT-3.5-Turbo. Notably, Qwen 2 (72B) is mostly on par with, sometimes superior to, GPT-4o (58.9-86.5% vs 67.3-86.9%). However, unlike GPT-4o, Qwen 2 (72B) generates longer summaries and has a higher hallucination rate (see Appendix E.2).
@@ 4.4 Output order: A Proxy for Correctness?
We observe that insights positioned earlier in the summary are more likely to be accurate than those located later (see Figure 5). In Appendix E.5, we show that this pattern holds across all summarization models, regardless of input document count. Interestingly, the opposite pattern appears in the top three summary positions. Manual inspection reveals that lower accuracy at the start is due to the generation of broader scope insights (e.g., opening remarks or preliminary statements). We also find that GPT-4o and Gemini (Flash) tend to generate 1 or 2 “takeaway” or “concluding” insights at the end of the summary to emphasize the main point or lesson intended to be remembered. We hypothesize this structure may stem from discourse coherence (Jurafsky and Martin, 2024; Zhu et al., 2024) or the models’ ability to mimic input document patterns. Future research could explore these possibilities further.
Finally, we also investigate which documents errors stem from, and find that most LLMs show a near-uniform distribution of hallucinations across input documents, while GPT-3.5-Turbo, Llama 3.1 (70B) and Gemini (Flash) display a slight recency bias towards the last documents. For detailed results of this experiment see Appendix E.6.
Figure 5: Accuracy rate of GPT-4o-generated insights by position (when summarizing 10 input documents). Each solid line shows the mean, with shaded areas representing 95% confidence intervals. Overall, accuracy rate declines as insight position increases.

### v2

@@ Abstract
Although many studies have investigated and reduced hallucinations in large language models (LLMs) for single-document tasks, research on hallucination in multi-document summarization (MDS) tasks remains largely unexplored. Specifically, it is unclear how the challenges arising from handling multiple documents (e.g., repetition and diversity of information) affect models outputs. In this work, we investigate how hallucinations manifest in LLMs when summarizing topic-specific information from a set of documents. Since no benchmarks exist for investigating hallucinations in MDS, we leverage existing news and conversation datasets, annotated with topic-specific insights, to create two novel multi-document benchmarks. When evaluating 5 LLMs on our benchmarks, we observe that on average, up to 75% of the content in LLM-generated summary is hallucinated, with hallucinations more likely to occur towards the end of the summaries. Moreover, when summarizing non-existent topic-related information, gpt-3.5-turbo and GPT-4o still generate summaries about 79.35% and 44% of the time, raising concerns about their tendency to fabricate content. To better understand the characteristics of these hallucinations, we conduct a human evaluation of 700+ insights and discover that most errors stem from either failing to follow instructions or producing overly generic insights. Motivated by these observations, we investigate the efficacy of simple post-hoc baselines in mitigating hallucinations but find them only moderately effective. Our results underscore the need for more effective approaches to systematically mitigate hallucinations in MDS.
@@
Catarina G. Belemthanks:  Work done while interning at Megagon Labs. University of California Irvine cbelem@uci.edu                        Pouya Pezeshkpour Megagon Labs pouya@megagon.ai                        Hayate Iso Megagon Labs hayate@megagon.ai
@@ 1 Introduction
Multi-document summarization (MDS) has numerous real-world applications, including planning treatments and diagnosing patients based on their medical history (i.e., doctor notes, lab reports) (Tang et al., 2023b), forming legal arguments by linking precedents (Rodgers et al., 2023; Wu et al., 2023), or screening resumes to match candidates to job descriptions (Wang et al., 2024; Du et al., 2024). These tasks often involve linking information across lengthy documents, making the summarization process time- and labor-intensive (Van Veen et al., 2024). Recently, large language models (LLMs) have been proposed as more efficient approaches to reduce the human effort and increase scalability (Katz et al., 2023; Van Veen et al., 2024; Liu et al., 2024).
@@
However, despite the remarkable advances of LLMs in various tasks (Bubeck et al., 2023; Zhao et al., 2023), the frequent generation of ungrounded yet plausible-sounding text, referred to as “hallucinations”, undermines trust in their output (Maynez et al., 2020; Uluoglakci and Temizel, 2024; Kalai and Vempala, 2024). While LLM hallucinations have been extensively studied in single-document tasks (Ji et al., 2023; Huang et al., 2025), resulting in the development of various evaluation benchmarks (Lin et al., 2022; Wang et al., 2022; Yang et al., 2023) and taxonomies (Rawte et al., 2023; Zhang et al., 2023; Mishra et al., 2024; Dahl et al., 2024), little is known about how processing multiple documents affects the hallucinatory behavior of LLMs in MDS. Specifically, by focusing on the aggregated quality metrics of LLM outputs, including coverage (Fabbri et al., 2019; Lu et al., 2020; Laban et al., 2024) and faithfulness (Pu et al., 2023; Huang et al., 2024), existing work offers limited insight into how the inherent challenges of MDS (e.g., repeated and diversity of information) correlate with LLM hallucinations. Consider the motivating example in Figure 1: when summarizing the repeated information from multiple documents, the LLM fails to link overlapping information across documents and satisfy the subtopic condition (highlighted text in the figure).
In this work, we comprehensively investigate the hallucinatory behavior of 5 popular LLMs in MDS tasks. Focusing on news articles and dialogues summarization, we examine the properties of hallucinations produced by LLMs when summarizing topic-specific information from multiple documents. Specifically, we aim to understand the patterns of hallucinations, their frequency, and their relationship with the number of input documents, as well as how these patterns change with the number of input documents and task focus (e.g., summarizing undiscussed topics, or focusing the summary on repeated information). To this end, we design an evaluation protocol based on fine-grained annotations concerning relevant insights—units of information—within each document. We use the insight-level annotations from the SummHay datasets (Laban et al., 2024), designed for evaluating LLMs in long-context summarization, and create benchmarks with combinations of up to 10 documents in both the conversational and news domains. Our goal here is to leverage the insight-level annotations to automatically assess the correctness of the LLM-generated summaries. [^1] The code and data have been made publicly available: https://github.com/megagonlabs/Hallucination_MDS.
@@ 2.1 Problem Formulation
We frame the MDS task as follows: given $N$ documents ($d_{1},\dots,d_{N}$), and $K$ conditions ($c_{1},...,c_{K}$), an LLM must generate a summary $\hat{\mathbf{y}}$ such that it satisfies all $K$ conditions and is grounded in the documents. Examples of conditions include matching specific topics, adhering to length constraints, or following particular writing styles. For simplicity, we assume that the $N$ documents in our formulation are of the same type (e.g., all documents are news articles), rather than from diverse sources (e.g., job descriptions and resumes, doctor notes and prescriptions). The summarization of multiple documents may involve capturing information that is either diverse (Huang et al., 2024), contradictory, or common across documents. The latter is especially important for combating misinformation, as it emphasizes “trustworthy” information supported by multiple sources. To analyze model behavior in both contexts, we require fine-grained annotations of the information units contained in each document.
@@ 2.2 Dataset Creation
Recently, Laban et al. (2024) introduced two MDS datasets with insight-level annotations. [^2] The authors define insights as “units of information”. The datasets cover two distinct domains—news and conversation—each containing 500 documents, evenly distributed across 5 topics, with documents averaging around 750 words. In particular, the news dataset (SummHay-News) is entity-centric and quantitative, featuring brands (e.g., Tesla), banks (e.g., JP Morgan), celebrities (e.g., Elon Musk), along with numbers and dates. In contrast, the conversation dataset (SummHay-Conv) involves everyday scenarios with 2 to 4 participants, such as medical appointments and debates. Originally developed to evaluate LLMs in long-context tasks (e.g., retriever-augmented generation, multi-document summarization), the documents were designed to ensure insights are repeated across at least 6 per topic and categorized into various subtopics, with no contradictory insights present.
By selecting SummHay-News and SummHay-Conv as testbeds, we can exploit their insight-level annotations to systematically assess model behavior across different settings, such as number of input documents and task focus. To control for input length (Li et al., 2024) while evaluating model behavior, we organize documents in the SummHay dataset into sets of $N$ documents, referred to as combinations of $N$ or $N$-documents (see Figure 7 in Appendix for an illustration of our benchmark creation process). Specifically, given a corpus with insight-level annotations and a subtopic $q$, we create our benchmarks by grouping documents into combinations of $N$ where multiple subtopic-related insights co-occur in two or more documents. Because we have the insight-level annotations, we can automatically identify the ground truth insights related to subtopic $q$ (called reference insights). To fully benefit from these fine-grained annotations, we instruct LLMs to succinctly333 Unlike Laban et al., we do not instruct LLMs to limit their generation to a fixed number of bullet-points. Instead, we study their behavior “in the wild” when the exact number of insights is unknown. For further details related to prompts, refer to Appendix B. summarize the information as individual insights, presented in the form of bullet-point lists. We coin this setting “subtopic”. To further investigate how the model handles common information, we introduce a “subtopic+trustworthy” setting by refining the prompt and restricting the reference insights to only those that are shared across documents.
@@
Note that while the proposed methodology applies to any MDS dataset with high-quality insight-level annotations, such information is rarely available. Therefore, we limit our evaluation to the SummHay dataset and leave the evaluation of other datasets, including non-English datasets and those with contradictory insights, for future work.
@@ 2.3 Automatic Evaluation
With reference insights for each combination, we can automatically evaluate the correctness of LLM outputs, distinguishing between hallucinatory and non-hallucinatory content. The correctness metric should assess whether predicted insights capture all essential information from the reference insights while avoiding the inclusion of irrelevant details (Huang et al., 2024). This ensures that predicted insights do not contain unsupported or potentially fabricated information.
Previous work uses an LLM-as-a-judge approach (Zheng et al., 2023) to assess whether reference insights are fully, partially, or not covered by the predicted insights, showing strong correlation with human evaluations (Laban et al., 2024). However, this metric only measures reference coverage and ignores the validity of additional details in predictions. To enhance this, we assess how well predicted insights align with reference insights by applying the original metric twice with swapped inputs—once for reference insight and once for predicted insight—–yielding two coverage labels per pair. We then combine these into a single correctness label using a conservative approach based on the order: NO $\prec$ PARTIAL $\prec$ FULL. By selecting the worst label (i.e., select NO over PARTIAL and select PARTIAL over FULL), we ensure the final assessment captures discrepancies related to missing relevant information or any potentially unfaithful content that is generated by the LLM.
@@ 2.4 Metrics
A key question in MDS tasks is whether LLM-generated summaries include all the correct information. To estimate average LLM correctness, we calculate macro-recall by determining the fraction of reference insights covered in each generated summary and averaging these scores across all summaries to obtain a single score. We also report the average fraction of hallucinated content in generated summaries by computing the average proportion of predicted insights that do not correspond to any reference insight across all summaries. This metric is referred to as the false discovery rate (macro-FDR). Both metrics range from 0 to 1, with a perfect LLM scoring 1 for macro-recall and 0 for macro-FDR. For simplicity, we refer to macro-recall as recall and macro-FDR as hallucination rate.
@@ 3 Experimental Settings
Automatic Evaluation. As previously mentioned, we adopt a few-shot LLM-as-a-judge approach. However, instead of using gpt-4o like Laban et al. (2024), we resort to a competitive yet more budget-friendly option—gpt-4o-mini-2024-07-18. The authors manually annotated 100 insight pairs and found strong alignment between both models, confirming the suitability of gpt-4o-mini-2024-07-18 as an evaluator. Refer to Appendix D for more details.
@@ 4 Experimental Results
In this section, we first investigate the extent of LLMs hallucinations in MDS under the proposed “subtopic” and “subtopic+trustworthy” settings. Then, we analyze LLMs’ hallucinatory behavior when synthesizing information on non-existent topics and investigate the link between hallucinated content and characteristics of both the input and output (we use the prompts listed in Appendix B).
@@
Figure 3: Performance metrics as a function of the number of input documents in the “subtopic+trustworthy” setting. Each line represents the mean value, with shaded areas indicating the 95% confidence intervals.
@@ 4.2 Shared Insights, Greater Errors
Compared to the previous “subtopic” setting, analyzing model behavior in “subtopic+trustworthy” requires adjusting the evaluation process and summarization prompt to focus only on shared insights. As before, we prompt models to summarize all 500 examples per benchmark and measure recall and hallucination rates. Our findings are as follows:
LLMs generate shorter summaries but hallucinate more in general. To understand how responsive models are to the shared instruction, we compare summary lengths between “subtopic” and “subtopic+trustworthy” settings (see Appendix E.2). Summaries in “subtopic+trustworthy” are shorter, suggesting models respond to the shared instruction. However, shorter summaries do not guarantee quality: hallucination rates are, on average, higher in “subtopic+trustworthy” than “subtopic” (news: +10.47%, conv: +4.20%), indicating models may struggle to identify shared insights. We also observe that, LLMs struggle to identify subtopic-related shared insights. Compared to the “subtopic” scenario, we observe an average increase in summarization performance in news domain (+6.93%) but a slight decrease in the conversation (-2.91%). Overall recall trends remain the same: larger combination sizes lead to up to a 33% drop, with the sharpest declines in GPT-3.5-Turbo. Notably, Qwen 2 (72B) is mostly on par with, sometimes superior to, GPT-4o (58.9-86.5% vs 67.3-86.9%). However, unlike GPT-4o, Qwen 2 (72B) generates longer summaries and has a higher hallucination rate (see Appendix E.2).
@@ 4.3 Summarizing the Unsummarizable
4.4 Which documents do errors stem from?
So far, we have presented empirical evidence that models consistently hallucinate, irrespective of the number of input documents or task focus. In this section, we investigate the source of model hallucinations by examining how they correlate with the input documents, i.e., we ask the question where do hallucinations come from? [^5] Prior work has studied this question with respect the faithfulness score (Huang et al., 2024). We tackle this question by first determining which input document (if any) the hallucinated insight is copied from. In particular, we modify the string matching approach used to determine data contamination in the GPT-4 paper (OpenAI, 2024b), leaving the exploration of other techniques (Xu et al., 2024) for future work. Specifically, we assume that an hallucinated insight $i$ originates from a document $d$ if any 50-character string from $i$ matches a string in $d$.
We find that 3 out of 5 models exhibit a slight recency bias towards the last documents. As observed in Figure 5, this recency bias is particularly prominent in GPT-3.5-Turbo and Gemini (Flash) in the news domain, but also present in GPT-3.5-Turbo and Llama 3.1 (70B) in the conversation domain. We hypothesize that such biases may emerge from the proximity to the list of instructions or due to the models’ positional encoding, whose exploration we leave for future work.
Figure 5: Likelihood of an insight coming from a document (y-axis) based on its position in the input (when summarizing 10 documents). GPT-3.5-Turbo, Llama 3.1 (70B) and Gemini (Flash) mistakes seem to be more likely to originate from later documents on average than from earlier ones.
@@ 4.5 Output order: A Proxy for Correctness?
We observe that insights positioned earlier in the summary are more likely to be accurate than those located later (see Figure 6), consistent with previous findings (Min et al., 2023; Chen et al., 2023a). In Appendix E.5, we show that this pattern holds across all summarization models, regardless of input document count. Interestingly, the opposite pattern appears in the top three summary positions. Manual inspection reveals that lower accuracy at the start is due to the generation of broader scope insights (e.g., opening remarks or preliminary statements). We also find that GPT-4o and Gemini (Flash) tend to generate 1 or 2 “takeaway” or “concluding” insights at the end of the summary to emphasize the main point or lesson to remember. We hypothesize this structure may stem from discourse coherence (Jurafsky and Martin, 2024; Zhu et al., 2024) or the models’ ability to mimic input document patterns. Future research could explore these possibilities further.
Figure 6: Accuracy rate of GPT-4o generated insights by position (when summarizing 10 input documents). Each solid line shows the mean, with shaded areas representing 95% confidence intervals. Overall, accuracy rate declines as insight position increases.